2.)Force necessary to pull a circular plate of 5cm radius from water surface for which surface tension is 75 dynes/cm, is
3.)A square frame of side L is dipped in a liquid. On taking it out, a membrane is formed. If the surface tension of the liquid is T, the force acting on the frame will be
4.)The material of a wire has a density of 1.4 g per cm3. If it is not wetted by a liquid of surface tension 44 dyne per cm, then the
maximum radius of the wire which can float on the surface of the liquid is

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
Please ask one question in one thread.
Answer of 1st question
F = T * I = 75 * 2 \pi * r = 150 * 5 \pi 
= 750 \pi dynes
Hope it helps
 
Regards
Arun (askIITians forum expert)